(s 2/3 power x t 1/2 power)(s 1/2 power x t 1/5 power) =
Answer by stanbon(48558)

(s 2/3 power x t 1/2 power)(s 1/2 power x t 1/5 power) =
(s^(2/3)t^(1/2))(s^(1/2)t^(1/5))
= s^(2/3+1/2)t^(1/2+1/5)
= s^(7/6)t^(7/10)
=========
Cheers,
Stan H.